# WavReward: Spoken Dialogue Models With Generalist Reward Evaluators

### v1

@@ Abstract
End-to-end spoken dialogue models such as GPT-4o-audio have recently garnered significant attention in the speech domain. However, the evaluation of spoken dialogue models’ conversational performance has largely been overlooked. This is primarily due to the intelligent chatbots convey a wealth of non-textual information which cannot be easily measured using text-based language models like ChatGPT. To address this gap, we propose WavReward, a reward feedback model based on audio language models that can evaluate both the IQ and EQ of spoken dialogue systems with speech input. Specifically, 1) based on audio language models, WavReward incorporates the deep reasoning process and the nonlinear reward mechanism for post-training. By utilizing multi-sample feedback via the reinforcement learning algorithm, we construct a specialized evaluator tailored to spoken dialogue models. 2) We introduce ChatReward-30K, a preference dataset used to train WavReward. ChatReward-30K includes both comprehension and generation aspects of spoken dialogue models. These scenarios span various tasks, such as text-based chats, nine acoustic attributes of instruction chats, and implicit chats. WavReward outperforms previous state-of-the-art evaluation models across multiple spoken dialogue scenarios, achieving a substantial improvement about Qwen2.5-Omni in objective accuracy from 55.1 to 91.5. In subjective A/B testing, WavReward also leads by a margin of 83. Comprehensive ablation studies confirm the necessity of each component of WavReward. All data and code will be publicly at https://github.com/jishengpeng/WavReward after the paper is accepted.
@@ 3.2.1 The Overall of ChatReward-30K
[…]
@@ 4.1 Experiment setup
Training details. WavReward is trained using 8 NVIDIA H20 96GB GPUs, each running a batch size of 1, with gradient accumulation performed every 2 steps. The model is trained for 3500 steps with a learning rate of and a temperature of 1.0. The maximum number of cot tokens is set to 5120, and the weight coefficient for the KL loss is set to 0.01. The model architecture of WavReward is based on the sota open-source audio language model Qwen2.5-Omni-7B-Think [25] with the identical parameters. All parameters of WavReward are updated during the training process.
@@ 4.2 Main Results
Model Content Acoustic Instruction Implicit RealDialogue age. acc. lan. gen. emo. pit. spe. vol. aud. I. Baseline audio langauge models direct inference with prompt Qwen2-Audio 24.7 32.4 24.5 36.8 27.6 33.7 32.3 40.5 41.4 50.8 28.9 42.5 Qwen-Audio 43.4 35.5 23.0 33.6 14.9 34.2 27.7 35.4 39.0 32.2 35.2 40.8 SALMONN 13.5 33.9 34.8 28.4 33.9 36.4 25.4 30.3 28.0 51.3 20.3 19.2 Audio Flamingo2 22.7 20.8 25.4 16.8 18.8 18.6 17.8 21.5 21.9 20.6 22.6 21.6 GPT-4o-audio 69.4 92.0 57.7 100 82.1 58.5 88.7 94.5 88.1 83.3 53.6 57.6 Qwen2.5-Omni 75.1 56.3 50.0 48.4 54.1 57.8 66.1 34.1 53.6 64.9 48.5 51.7 II. Baseline audio langauge models after supervised fine-tuning Qwen2.5-Omni w/ Full-param tuning 37.4 61.4 69.6 64.2 67.1 66.5 59.8 57.0 74.4 54.6 42.2 37.5 Qwen2.5-Omni w/ Lora 63.8 81.2 43.6 100 82.1 49.1 74.1 83.6 85.1 85.7 54.2 56.7 III. Different ablation versions of WavReward WavReward w/o cot think 84.2 80.3 77.9 98.9 86.9 85.4 80.7 86.0 90.2 88.6 61.4 59.1 WavReward w/o multi samples 85.3 85.7 69.6 98.9 88.6 90.0 82.1 88.6 85.4 90.2 61.9 72.5 WavReward w/o nonlinear reward 88.6 92.2 80.9 100 89.8 90.5 83.8 87.3 92.7 85.7 66.6 70.8 WavReward (ours) 90.8 96.9 87.7 100 95.5 97.5 89.1 91.1 97.6 97.0 74.3 80.8
@@
We evaluated the generation and ground-truth score accuracy of WavReward and Baseline models on the ChatReward-30K-test as well as the real out-of-domain RealDialogue dataset. The Baseline is divided into two categories: one consists of direct inference from audio language models using text prompt templates consistent with WavReward, and the other is the evaluator fine-tuned using the ChatReward-30K-train. The specific experimental results are presented in Table 2. We can draw the following conclusions: 1) WavReward significantly outperforms the best audio language models GPT-4o-audio on all metrics. It achieved improvements of 21.4, 20.7, and 39.0 points in the content scoring, implicit dialogue scoring, and emotion-instructed dialogue scoring, respectively. Furthermore, it outperformed the direct inference model Qwen2.5-Omni by an average factor of two. This indicates that audio language models when optimized using reinforcement learning, can effectively serve as evaluators for spoken dialogue models. Moreover, RL significantly improves performance compared to direct inference. 2) We found that the RL-based WavReward surpassed the LoRA fine-tuned Qwen2.5-Omni. This may be due to the direct scoring approach of supervised fine-tuning, which is overly simplistic and struggles to capture the complex scoring logic needed for various scenarios. Additionally, LoRA-based methods outperformed full-parameter fine-tuning, likely because full-parameter fine-tuning lacks the KL loss constraint on the initial model, causing the model to deviate from its inherent capabilities. 3) We observed a substantial performance gap between different audio language models during direct inference, highlighting the need for future work to develop and open-source more robust foundational audio language models. 4) We found that WavReward’s accuracy in scoring accent-based sub-languages and implicit dialogues was lower than that for other scenarios. This may be attributed to the less accurate accent data in the ChatReward-30K dataset compared to other acoustic information, and the inherent difficulty in evaluating implicit dialogues. There remains room for improvement in determining what constitutes a reasonable emotional response in spoken dialogue models. 5) On the RealDialogue dataset, WavReward achieved a score accuracy of 80, indicating that it exhibits strong robustness and can provide reliable evaluations in real-world, complex scenarios.

### v2

@@ Abstract
End-to-end spoken dialogue models such as GPT-4o-audio have recently garnered significant attention in the speech domain. However, the evaluation of spoken dialogue models’ conversational performance has largely been overlooked. This is primarily due to the intelligent chatbots convey a wealth of non-textual information which cannot be easily measured using text-based language models like ChatGPT. To address this gap, we propose WavReward, a reward feedback model based on audio language models that can evaluate both the IQ and EQ of spoken dialogue systems with speech input. Specifically, 1) based on audio language models, WavReward incorporates the deep reasoning process and the nonlinear reward mechanism for post-training. By utilizing multi-sample feedback via the reinforcement learning algorithm, we construct a specialized evaluator tailored to spoken dialogue models. 2) We introduce ChatReward-30K, a preference dataset used to train WavReward. ChatReward-30K includes both comprehension and generation aspects of spoken dialogue models. These scenarios span various tasks, such as text-based chats, nine acoustic attributes of instruction chats, and implicit chats. WavReward outperforms previous state-of-the-art evaluation models across multiple spoken dialogue scenarios, achieving a substantial improvement about Qwen2.5-Omni in objective accuracy from 53.4 to 91.5. In subjective A/B testing, WavReward also leads by a margin of 83. Comprehensive ablation studies confirm the necessity of each component of WavReward. All data and code will be publicly at https://github.com/jishengpeng/WavReward after the paper is accepted.
@@ 3.2.1 The Overall of ChatReward-30K
Dataset/Benchmark Dia. S2S. Neg. Imp. Sco. Acoustic Paralinguistic Information age. acc. lan. gen. emo. pit. spe. vol. aud. SUPERB MMAU AudioBench AirBench SD-Eval VoiceBench VoxDialogue ChatReward-30K
@@ 4.1 Experiment setup
Training details. WavReward is trained using 8 GPUs (due to company policy, the specific details are confidential.), each running a batch size of 1, with gradient accumulation performed every 2 steps. The model is trained for 3500 steps with a learning rate of and a temperature of 1.0. The maximum number of cot tokens is set to 5120, and the weight coefficient for the KL loss is set to 0.01. The model architecture of WavReward is based on the sota open-source audio language model Qwen2.5-Omni-7B-Think [25] with the identical parameters. All parameters of WavReward are updated during the training process.
@@ 4.2 Main Results
Model Content Acoustic Instruction Implicit RealDialogue age. acc. lan. gen. emo. pit. spe. vol. aud. I. Baseline audio langauge models direct inference with prompt Qwen2-Audio 24.7 32.4 24.5 36.8 27.6 33.7 32.3 40.5 41.4 50.8 28.9 42.5 Qwen-Audio 43.4 35.5 23.0 33.6 14.9 34.2 27.7 35.4 39.0 32.2 35.2 40.8 SALMONN 13.5 33.9 34.8 28.4 33.9 36.4 25.4 30.3 28.0 51.3 20.3 19.2 Audio Flamingo2 22.7 20.8 25.4 16.8 18.8 18.6 17.8 21.5 21.9 20.6 22.6 21.6 GPT-4o-audio 69.4 92.0 57.7 100 82.1 58.5 88.7 94.5 88.1 83.3 53.6 57.6 Qwen2.5-Omni 54.6 56.3 50.0 48.4 54.1 57.8 66.1 34.1 53.6 64.9 48.5 51.7 II. Baseline audio langauge models after supervised fine-tuning Qwen2.5-Omni w/ Full-param tuning 67.1 81.9 69.6 98.9 83.8 66.5 76.1 84.8 86.6 86.7 55.8 58.3 Qwen2.5-Omni w/ LoRA 63.8 81.2 43.6 100 82.1 49.1 74.1 83.6 85.1 85.7 54.2 56.7 III. Different ablation versions of WavReward WavReward w/o cot think 84.2 80.3 77.9 98.9 86.9 85.4 80.7 86.0 90.2 88.6 61.4 59.1 WavReward w/o multi samples 85.3 85.7 69.6 98.9 88.6 90.0 82.1 88.6 85.4 90.2 61.9 72.5 WavReward w/o nonlinear reward 88.6 92.2 80.9 100 89.8 90.5 83.8 87.3 92.7 85.7 66.6 70.8 WavReward (ours) 90.8 96.9 87.7 100 95.5 97.5 89.1 91.1 97.6 97.0 74.3 80.8
@@
We evaluated the generation and ground-truth score accuracy of WavReward and Baseline models on the ChatReward-30K-test as well as the real out-of-domain RealDialogue dataset. The Baseline is divided into two categories: one consists of direct inference from audio language models using text prompt templates consistent with WavReward, and the other is the evaluator fine-tuned using the ChatReward-30K-train. The specific experimental results are presented in Table 2. We can draw the following conclusions: 1) WavReward significantly outperforms the best audio language models GPT-4o-audio on all metrics. It achieved improvements of 21.4, 20.7, and 39.0 points in the content scoring, implicit dialogue scoring, and emotion-instructed dialogue scoring, respectively. Furthermore, it outperformed the direct inference model Qwen2.5-Omni by an average factor of two. This indicates that audio language models when optimized using reinforcement learning, can effectively serve as evaluators for spoken dialogue models. Moreover, RL significantly improves performance compared to direct inference. 2) We found that the RL-based WavReward surpassed the LoRA fine-tuned Qwen2.5-Omni. This may be due to the direct scoring approach of supervised fine-tuning, which is overly simplistic and struggles to capture the complex scoring logic needed for various scenarios. 3) We observed a substantial performance gap between different audio language models during direct inference, highlighting the need for future work to develop and open-source more robust foundational audio language models. 4) We found that WavReward’s accuracy in scoring accent-based sub-languages and implicit dialogues was lower than that for other scenarios. This may be attributed to the less accurate accent data in the ChatReward-30K dataset compared to other acoustic information, and the inherent difficulty in evaluating implicit dialogues. There remains room for improvement in determining what constitutes a reasonable emotional response in spoken dialogue models. 5) On the RealDialogue dataset, WavReward achieved a score accuracy of 80, indicating that it exhibits strong robustness and can provide reliable evaluations in real-world, complex scenarios.